OpenStudy (anonymous):

Evaluate the expression (5+10i)(6-8i) and write your answer in the form a+bi where a,b are real numbers.

OpenStudy (anonymous):

(5+10i)(6-8i)=30-40i+60i-80i^2 i^2=-1 so 110+20i

OpenStudy (anonymous):

Thanks thats what I ended up getting! :)